# Resource-Optimized Grouping Shadow for Efficient Energy Estimation

### v1

@@ 1 Introduction
In this work, we propose Resource-Optimized Grouping Shadow (ROGS), a novel approach for the efficient estimation of Pauli observables’ expectation values on quantum states. Similar to the previous works [15, 16, 17, 18, 19, 20], ROGS also acts as a pre-processing stage for the classical shadow protocol [^1] Although we refer to our method as a shadow method, the techniques introduced here have potential applications beyond classical shadows. See more discussions in Section 6.. We consider efficiency from two aspects: the number of measurements ($n_{\mathrm{shot}}$) and the number of circuits ($n_{\mathrm{circuit}}$) prepared for measurements. In the context of this paper, a circuit refers to a specific sequence of quantum gates applied to qubits, followed by measurements in the computational basis (the $z$-basis) [^2] The measurement circuits consist solely of single-qubit gates, specifically Pauli measurement gates. Adopting the notation of [14], each Pauli measurement gate is composed of a single-qubit rotational gate followed by a measurement in the computational basis (z-basis). The rotational gate of the circuit is the tensor product of unitaries on each qubit, $U=\bigotimes_{i=1}^{n}U_{i}$, where $U_{i}$ is selected from the set $\{H,S^{\dagger}H,I\}$. Here, $H$ denotes the Hadamard gate, $S^{\dagger}$ is the adjoint of the phase gate, and $I$ is the identity gate. This combination of local rotational gates and z-basis measurements effectively realizes Pauli-basis measurements, enabling the estimation of expectation values for Pauli observables. . A shot or measurement refers to a single execution of this circuit on the quantum hardware, yielding a bitstring ($b\in\{\pm 1\}^{\otimes n}$) outcome. When estimating non-commuting observables, we need to measure multiple distinct circuits, each corresponding to a different Pauli operator. Here we refer to the number of measurements as the number of total shots we are allowed to gather from the quantum device or the number of times the device is queried; the measurement circuits are referred to as the distinct circuits used in the measurement process.
@@ 3.1 Max-Min Grouping Protocol
    $\{\mathcal{C}_{\alpha}\}=\left\{\arg\min_{\{\mathcal{C}_{\alpha}\}}|\{\mathcal{C}_{\alpha}\}\Bigg{\lvert}[O_{i},O_{j}]_{\mathrm{QWC}}=0,\leavevmode\nobreak\ \forall O_{i},O_{j}\in\mathcal{C}_{\alpha},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cup_{\alpha}\mathcal{C}_{\alpha}=\{O_{\ell}\}_{\ell\leq L},\leavevmode\nobreak\ \mathcal{C}_{\alpha}\cap\mathcal{C}_{\beta}=\emptyset,\leavevmode\nobreak\ \forall\alpha\neq\beta\right\}$ (3)
@@
Figure 1: A graphical illustration of our Max-Min grouping protocol. Each vertex represents a Pauli observable, with individual Pauli operators denoted by squares of varying colors. Left [Equation 3: Minimize]: solid lines connect vertices that commute with each other. One of the minimum clique covers is depicted, with operators within the same shadows belonging to the same group $(\mathcal{C}_{\alpha})_{\alpha=1,2,3}$. Middle [Equation 4: Maximize]: the process of maximizing overlaps between groups to obtain $\mathcal{C}^{\mathrm{MM}}_{\alpha}$ is shown. In this example, operators inside the solid gray line are all operators that commute with $\mathcal{C}_{2}$, and the maximum clique of those operators that commute with each other (inside the gray dashed line) is the $\mathcal{C}_{2}^{\mathrm{add}}$ we are looking for to include into $\mathcal{C}_{2}$, and so it is with $\mathcal{C}_{3}^{\mathrm{add}}$ (circled by the yellow dashed line). Right [Equation 5]: the overlapped shaded region $\mathcal{C}_{2}^{\mathrm{add}}$ includes operators in $\mathcal{C}_{1}$ that also commute with all operators in $\mathcal{C}_{1}^{\mathrm{MM}}=\mathcal{C}_{1}$ and $\mathcal{C}_{2,3}^{\mathrm{MM}}=\mathcal{C}_{2,3}\cup\mathcal{C}_{2,3}^{\mathrm{add}}$. The Pauli gates associated with group $\mathcal{C}_{1,2,3}$ are $\mathcal{P}_{1}=X_{1}\otimes X_{2}\otimes Z_{3}$, $\mathcal{P}_{2}=X_{1}\otimes Z_{2}\otimes I_{3}$ and $\mathcal{P}_{3}=X_{1}\otimes Y_{2}\otimes X_{3}$; however, with the maximization procedure considered, measurement gates for $\mathcal{C}^{\mathrm{MM}}_{2}$ will have a fixed measurement gate for the third qubit and changes to $\mathcal{P}_{2}=X_{1}\otimes Z_{2}\otimes Z_{3}$.
@@
The maximize process mainly helps us decide the measurement gates of qubits that have not been assigned by the minimize procedure. For example, in Fig. 1, the Pauli gates associated with $\mathcal{C}_{2}$ is changed from $X_{1}\otimes Z_{2}\otimes I_{3}$ to $X_{1}\otimes Z_{2}\otimes Z_{3}$ upon including $\mathcal{C}_{2}^{\mathrm{add}}$. This max-min grouping procedure ensures full utilization of each measurement. This maximization step is realized in Section 3.1 through steps $3$ to $12$. When we perform measurement for the group $\alpha$, using the measurement gates $\mathcal{P}_{\alpha}$, an operator will be measured if it belongs to the group $\alpha$. The overlap maximization step in our approach can increase the hit rates of observables that belong to the overlaps between operator groups. By maximizing the overlap, we ensure that the observables are measured more frequently within a fixed measurement budget, thereby improving the overall estimation accuracy. In general, this procedure guarantees high hit rates on observables, which is crucial for achieving precise estimates of the target quantities. The next task is to allocate measurement resources to each group to minimize the error.
@@ 3.2 Confidence Bound
In most scenarios, the confidence bound provided in Theorem 1 offers an effective guide for optimizing the allocation of measurement resources across groups, prioritizing the overall accuracy of the Hamiltonian estimation. However, in certain extreme cases, such as when a Hamiltonian contains groups with significantly imbalanced sizes and coefficient magnitudes (see Section A.2.1). the confidence bound in Theorem 1 may not provide the optimal allocation [^6] Although Theorem 1 does not yield optimal bounds without sampling in the hyperparameter space, it will produce accurate values if we follow the process detailed in Appendix C.. In such cases, the supplementary confidence bound introduced in Section A.2 can be more useful. This alternative bound takes into account the coefficients of the observables within each group, allowing for a more nuanced allocation of measurement resources. By considering each group’s relative importance based on its size and the magnitudes of its observable coefficients, the supplementary confidence bound can help mitigate the potential breakdown of the primary bound in these extreme scenarios. Nonetheless, for most practical applications, the confidence bound in Theorem 1 remains a robust and reliable tool for guiding the measurement optimization process.
@@ 3.3 Measurement Resource Allocation
2A fixed Pauli basis measurement recipe $\{\mathcal{P}_{\alpha}^{\otimes M_{\alpha}}\}_{1\leq\alpha\leq A}\in\{X,Y,Z\}^{n\times M}$ for the classical shadow quantum measurement stage.
@@
5Obtain measurement recipes $\{\mathcal{P}_{\alpha}^{\otimes M_{\alpha}}\}_{1\leq\alpha\leq A}$ which include $\{X,Y,Z\}^{n\times M}$ based on $\{M_{\alpha}\}:=\{[w_{\alpha}^{*}M]\}$ and one to one correspondence between groups $\{\mathcal{C}_{\alpha}^{\mathrm{MM}}\}$ and Pauli measurement gates $\mathcal{P}_{\alpha}$.
Measurement Resource Allocation via Convex Optimization. Now, we can determine the allocation of the measurement budget among groups based on the confidence bound (1) derived in Theorem 1. Note that the confidence bound can be transformed into a log-sum-exp function, which is smooth and convex. Minimization of this function can be efficiently performed by a convex optimization solver (e.g., CVXPY [26]), which outputs a set of optimal weights $\{w_{\alpha}\}$ for shot allocation to each group. We demonstrate the whole process in Section 3.3, which takes a measurement budget $M$ and outputs measurement recipes777$\{\mathcal{P}^{(m)}_{\alpha}\}_{1\leq m\leq M_{\alpha}}$ denotes the repetition of $\mathcal{P}_{\alpha}$ for $M_{\alpha}$ times, where $\mathcal{P}_{\alpha}=P_{1}^{\alpha}\otimes\dots\otimes P_{n}^{\alpha}$ ($P_{i}$ is a Pauli measurement gate on site-$i$) and is associated with a single Pauli measurement on the given state, yielding a single classical snapshot [14] $\hat{\rho}_{\alpha}$. The subscript $\alpha$ labels this classical snapshot because all operators $O_{\ell}$ in ${\mathcal{H}}$ that belong to $\mathcal{C}_{\alpha}^{\mathrm{MM}}$ produce a non-zero random variable $\mathrm{Tr}(\hat{\rho}_{\alpha}O_{\ell})\in\{-1,1\}$, which is used in predicting the expectation $\hat{O}_{\ell}=\mathbb{E}[\mathrm{Tr}(\hat{\rho}_{\alpha}O_{\ell})]$. Conversely, if $O_{\ell}\notin\mathcal{C}_{\alpha}^{\mathrm{MM}}$, this snapshot cannot be used to evaluate the expectation i.e. $\mathrm{Tr}(\hat{\rho}_{\alpha}O_{\ell})=0$ for all $O_{\ell}\notin\mathcal{C}_{\alpha}^{\mathrm{MM}}$., $\{\{\mathcal{P}^{(m)}_{\alpha}\}_{1\leq m\leq M_{\alpha}}\}$.
@@ 5 Discussion
Physical Interpretation. from a physical standpoint: each group $\mathcal{C}_{\alpha}^{\mathrm{MM}}$ can be viewed as a mean-field subspace of the Hamiltonian [33] [^8] In the mean-field approach, the Hamiltonian is partitioned into a sum of mean-field Hamiltonians, each of which has eigenstates that are unentangled products of single-qubit states. This allows each mean-field Hamiltonian to be measured using a single set of single-qubit measurements in the appropriate single-qubit eigenbases. However, unlike in our context, the terms within each mean-field Hamiltonian do not necessarily commute with each other.. These subspaces are constructed by maximizing the number of mutually commuting operators within each group, essentially capturing the local mean-field interactions among the qubits. The mean-field approximation is often used to describe systems where the interactions between particles can be effectively averaged out, leading to a simplified description of the system. In the context of the overlapped grouping protocol (Sec. 3.1), the larger groups with more mutually commuting operators can be seen as capturing these averaged interactions, providing a good approximation of the overall system behavior.
@@ Acknowledgments
The authors thank Katharine Hyatt, Eric Kessler, Péter Kómár, Cedric Yen-Yu Lin, and Jiayu Shen for inspiring discussions and valuable input. Special thanks to Eric Kessler and Cedric Yen-Yu Lin for their meticulous review, valuable advice, and numerous revisions.

### v2

@@ 1 Introduction
In this work, we propose Resource-Optimized Grouping Shadow (ROGS), a novel approach for the efficient estimation of Pauli observables’ expectation values on quantum states. Similar to the previous works [15, 16, 17, 18, 19, 20], ROGS also acts as a pre-processing stage for the classical shadow protocol [^1] Although we refer to our method as a shadow method, the techniques introduced here have potential applications beyond classical shadows. See more discussions in Section 6.. We consider efficiency from two aspects: the number of measurements ($n_{\mathrm{shot}}$) and the number of circuits ($n_{\mathrm{circuit}}$) prepared for measurements. In the context of this paper, a circuit refers to a specific sequence of quantum gates applied to qubits, followed by measurements in the computational basis (the $z$-basis) [^2] The measurement circuits consist solely of single-qubit gates, specifically Pauli measurement gates. Adopting the notation of [14], each Pauli measurement gate is composed of a single-qubit rotational gate followed by a measurement in the computational basis (z-basis). The rotational gate of the circuit is the tensor product of unitaries on each qubit, $U=\bigotimes_{i=1}^{n}U_{i}$, where $U_{i}$ is selected from the set $\{H,S^{\dagger}H,I\}$. Here, $H$ denotes the Hadamard gate, $S^{\dagger}$ is the adjoint of the phase gate, and $I$ is the identity gate. This combination of local rotational gates and z-basis measurements effectively realizes Pauli-basis measurements, enabling the estimation of expectation values for Pauli observables. . A shot or measurement refers to a single execution of this circuit on the quantum hardware, yielding a bitstring ($b\in\{\pm 1\}^{\otimes n}$) outcome. When estimating non-QWC observables, we need to measure multiple distinct circuits, each corresponding to a different Pauli operator. Here we refer to the number of measurements as the number of total shots we are allowed to gather from the quantum device or the number of times the device is queried; the measurement circuits are referred to as the distinct circuits used in the measurement process.
@@ 3.1 Max-Min Grouping Protocol
    $\{\mathcal{C}_{\alpha}\}=\left\{\arg\min_{\{\mathcal{C}_{\alpha}\}}|\{\mathcal{C}_{\alpha}\}|\Bigg{\lvert}[O_{i},O_{j}]_{\mathrm{QWC}}=0,\leavevmode\nobreak\ \forall O_{i},O_{j}\in\mathcal{C}_{\alpha},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cup_{\alpha}\mathcal{C}_{\alpha}=\{O_{\ell}\}_{\ell\leq L},\leavevmode\nobreak\ \mathcal{C}_{\alpha}\cap\mathcal{C}_{\beta}=\emptyset,\leavevmode\nobreak\ \forall\alpha\neq\beta\right\}$ (3)
@@
Figure 1: A graphical illustration of our Max-Min grouping protocol. Each vertex, which is depicted as a Pauli string, represents a Pauli observable, with individual Pauli operators denoted by squares of varying colors. Left [Equation 3: Minimize]: solid lines connect vertices that commute with each other. One of the minimum clique covers is depicted, with operators within the same shadows belonging to the same group $(\mathcal{C}_{\alpha})_{\alpha=1,2,3}$. Middle [Equation 4: Maximize]: the process of maximizing overlaps between groups to obtain $\mathcal{C}^{\mathrm{MM}}_{\alpha}$ is shown. In this example, operators inside the solid gray line are all operators that commute with $\mathcal{C}_{2}$, and the maximum clique of those operators that commute with each other (inside the gray dashed line) is the $\mathcal{C}_{2}^{\mathrm{add}}$ we are looking for to include into $\mathcal{C}_{2}$, and so it is with $\mathcal{C}_{3}^{\mathrm{add}}$ (circled by the yellow dashed line). Right [Equation 5]: the overlapped shaded region $\mathcal{C}_{2}^{\mathrm{add}}$ includes operators in $\mathcal{C}_{1}$ that also commute with all operators in $\mathcal{C}_{1}^{\mathrm{MM}}=\mathcal{C}_{1}$ and $\mathcal{C}_{2,3}^{\mathrm{MM}}=\mathcal{C}_{2,3}\cup\mathcal{C}_{2,3}^{\mathrm{add}}$. The Pauli gates associated with group $\mathcal{C}_{1,2,3}$ are $\mathcal{P}_{1}=X_{1}\otimes X_{2}\otimes Z_{3}$, $\mathcal{P}_{2}=X_{1}\otimes Z_{2}\otimes I_{3}$ and $\mathcal{P}_{3}=X_{1}\otimes Y_{2}\otimes X_{3}$; however, with the maximization procedure considered, measurement gates for $\mathcal{C}^{\mathrm{MM}}_{2}$ will have a fixed measurement gate for the third qubit and changes to $\mathcal{P}_{2}=X_{1}\otimes Z_{2}\otimes Z_{3}$.
@@
The maximize process mainly helps us decide the measurement gates of qubits that have not been assigned by the minimize procedure. For example, in Fig. 1, the Pauli gates associated with $\mathcal{C}_{2}$ is changed from $X_{1}\otimes Z_{2}\otimes I_{3}$ to $X_{1}\otimes Z_{2}\otimes Z_{3}$ upon including $\mathcal{C}_{2}^{\mathrm{add}}$. This max-min grouping procedure ensures full utilization of each measurement. This maximization step is realized in Section 3.1 through steps $3$ to $12$. When we perform measurement for the group $\alpha$, using the measurement gates $\mathcal{P}_{\alpha}$, an operator will be measured if it belongs to the group $\alpha$. The overlap maximization step in our approach can increase the hit rates of observables that belong to the overlaps between operator groups. By maximizing the overlap, we ensure that the observables are measured more frequently within a fixed measurement budget, thereby improving the overall estimation accuracy. In general, this procedure guarantees high hit rates on observables, which is crucial for achieving precise estimates of the target quantities.
In the Overlapped Grouping Measurement (OGM) framework [19], as well as in other grouping schemes, the central task is to find compatible sets of Pauli operators for simultaneous measurement. While our Max–Min grouping strategy is guided by the Minimum Clique Cover (MCC) heuristic (see discussion above), OGM and other related methods (e.g., [20]) often adopt a simpler largest degree first (LDF) rule or other greedy approaches. Generally, LDF aims to reduce the problem complexity by iteratively placing observables with the highest graph degree (i.e., the most edges in a compatibility graph) into existing or newly created cliques. This scheme often provides a decent partition without fully solving the MCC problem (which is NP-hard), but it can be suboptimal in cases where local high-degree choices do not translate into overall optimal grouping. In contrast, heuristics inspired by MCC may identify larger compatible sets and thus yield fewer final groups in practice, depending on the specific algorithm used. On the other hand, LDF and similar degree-based algorithms tend to be faster to implement in large-scale settings, at the risk of increasing the number of overall groups. In practice, which approach is preferable often depends on the trade-off between computational overhead for determining groups and the downstream cost of executing distinct measurement circuits. Our MCC-based Max–Min grouping leverages the clique cover perspective to minimize the circuit count while still accommodating overlaps, thereby giving both a systematic criterion for forming QWC groups and—once expanded—maximizing overlap among them. This ensures efficient shot usage and avoids redundant basis measurements, which can be crucial in resource-limited scenarios.
The next task is to allocate measurement resources to each group to minimize the error.
@@ 3.2 Confidence Bound
In most scenarios, the confidence bound provided in Theorem 1 offers an effective guide for optimizing the allocation of measurement resources across groups, prioritizing the overall accuracy of the Hamiltonian estimation. However, in certain extreme cases, such as when a Hamiltonian contains groups with significantly imbalanced sizes and coefficient magnitudes (see Section A.2.1), the confidence bound in Theorem 1 may not provide the optimal allocation [^6] Although Theorem 1 does not yield optimal bounds without sampling in the hyperparameter space, it will produce accurate values if we follow the process detailed in Appendix C.. In such cases, the supplementary confidence bound introduced in Section A.2 can be more useful. This alternative bound takes into account the coefficients of the observables within each group, allowing for a more nuanced allocation of measurement resources. By considering each group’s relative importance based on its size and the magnitudes of its observable coefficients, the supplementary confidence bound can help mitigate the potential breakdown of the primary bound in these extreme scenarios. Nonetheless, for most practical applications, the confidence bound in Theorem 1 remains a robust and reliable tool for guiding the measurement optimization process.
@@ 3.3 Measurement Resource Allocation
2A fixed Pauli basis measurement recipe $\{\{\mathcal{P}^{(m)}_{\alpha}\}_{1\leq m\leq M_{\alpha}}\}\in\{X,Y,Z\}^{n\times M}$ for the classical shadow quantum measurement stage.
@@
5Obtain measurement recipes $\{\{\mathcal{P}^{(m)}_{\alpha}\}_{1\leq m\leq M_{\alpha}}\}$ which include $\{X,Y,Z\}^{n\times M}$ based on $\{M_{\alpha}\}:=\{[w_{\alpha}^{*}M]\}$ and one to one correspondence between groups $\{\mathcal{C}_{\alpha}^{\mathrm{MM}}\}$ and Pauli measurement gates $\{\mathcal{P}_{\alpha}\}$.
Now, we can determine the allocation of the measurement budget among groups based on the confidence bound (1) derived in Theorem 1. In essence, this bound provides a way to quantify how the total estimation error probability depends on the distribution of the measurements (shots) among the groups. By transforming the confidence bound into a log-sum-exp function (see Eq. (1)), we obtain a smooth and convex objective that naturally lends itself to efficient numerical minimization.
Concretely, we introduce continuous variables $\{w_{\alpha}\}$, each representing the fraction of the total measurement budget $M$ allocated to group $\mathcal{C}_{\alpha}^{\mathrm{MM}}$. We then solve eq. 11 where “$\mathrm{Conf}[\dots]$” refers to the log-sum-exp confidence bound and $\epsilon$ is an error parameter. The solution $\{w_{\alpha}^{\ast}\}$ provides an optimal relative shot allocation across groups, in the sense that it achieves the smallest upper bound on the probability of exceeding the target estimation error $\epsilon$. Since $\mathrm{Conf}[\dots]$ has a known convex form, we can use standard convex optimization toolboxes such as CVXPY [26] to solve for $\{w_{\alpha}^{\ast}\}$ efficiently, even for relatively large problems.
In effect, the solver balances the measurement ‘importance’ of each group. Groups with many high-coefficient or critical operators (in terms of their Hamiltonian contribution) reduce the overall error bound more significantly when they receive more shots. This phenomenon will be discussed with greater detail in Section 5. Once the optimal fractions $\{w_{\alpha}^{\ast}\}$ are obtained, we translate them directly into integer numbers of shots $M_{\alpha}=\lfloor w_{\alpha}^{\ast}\,M\rfloor$. Each $M_{\alpha}$ is then used to measure the group $\mathcal{C}_{\alpha}^{\mathrm{MM}}$ repeatedly under the same Pauli measurement circuit $\mathcal{P}_{\alpha}$. The set of all such circuits and shot counts $\{\{\mathcal{P}_{\alpha}^{(m)}\}_{1\leq m\leq M_{\alpha}}\}$, constitutes our measurement recipe777$\{\mathcal{P}^{(m)}_{\alpha}\}_{1\leq m\leq M_{\alpha}}$ denotes the repetition of $\mathcal{P}_{\alpha}$ for $M_{\alpha}$ times, where $\mathcal{P}_{\alpha}=P_{1}^{\alpha}\otimes\dots\otimes P_{n}^{\alpha}$ ($P_{i}$ is a Pauli measurement gate on site-$i$) and is associated with a single Pauli measurement on the given state, yielding a single classical snapshot [14] $\hat{\rho}_{\alpha}$. The subscript $\alpha$ labels this classical snapshot because all operators $O_{\ell}$ in ${\mathcal{H}}$ that belong to $\mathcal{C}_{\alpha}^{\mathrm{MM}}$ produce a non-zero random variable $\mathrm{Tr}(\hat{\rho}_{\alpha}O_{\ell})\in\{-1,1\}$, which is used in predicting the expectation $\hat{O}_{\ell}=\mathbb{E}[\mathrm{Tr}(\hat{\rho}_{\alpha}O_{\ell})]$. Conversely, if $O_{\ell}\notin\mathcal{C}_{\alpha}^{\mathrm{MM}}$, this snapshot cannot be used to evaluate the expectation i.e. $\mathrm{Tr}(\hat{\rho}_{\alpha}O_{\ell})=0$ for all $O_{\ell}\notin\mathcal{C}_{\alpha}^{\mathrm{MM}}$. for energy estimation. In this way, the total measurement budget $M$ is divided among distinct circuits to optimize the overall accuracy of our Hamiltonian estimator.
@@ 5 Discussion
Let us comment on the physical interpretation of the ROGS method. From a physical standpoint, each group $\mathcal{C}_{\alpha}^{\mathrm{MM}}$ can be viewed as a mean-field subspace of the Hamiltonian [33] [^8] In the mean-field approach, the Hamiltonian is partitioned into a sum of mean-field Hamiltonians, each of which has eigenstates that are unentangled products of single-qubit states. This allows each mean-field Hamiltonian to be measured using a single set of single-qubit measurements in the appropriate single-qubit eigenbases. However, unlike in our context, the terms within each mean-field Hamiltonian do not necessarily commute with each other.. These subspaces are constructed by maximizing the number of mutually commuting operators within each group, essentially capturing the local mean-field interactions among the qubits. The mean-field approximation is often used to describe systems where the interactions between particles can be effectively averaged out, leading to a simplified description of the system. In the context of the overlapped grouping protocol (Sec. 3.1), the larger groups with more mutually commuting operators can be seen as capturing these averaged interactions, providing a good approximation of the overall system behavior.
@@ 6 Conclusion and Outlook
While our protocol substantially reduces the number of measurement circuits and hence some classical overhead, it does not inherently account for hardware noise that can arise in NISQ devices. In particular, noise and gate imperfections can distort measurement outcomes in the low-shot regime, which is precisely the scenario where adaptive shot-allocation methods (see Appendix C) are most sensitive. Consequently, one may need to incorporate additional error mitigation tools — such as zero-noise extrapolation or probabilistic error cancellation — to preserve the fidelity of the allocated measurements. These methods can be integrated into our grouping-and-allocation framework so that the benefits of reduced circuit count and targeted measurement remain effective despite device-level imperfections.
@@ Acknowledgments
The authors thank Katharine Hyatt, Eric Kessler, Péter Kómár, Cedric Yen-Yu Lin, and Jiayu Shen for inspiring discussions and valuable input. Special thanks to Eric Kessler for his meticulous review, valuable advice, and numerous revisions.